# Dual-Stage Cross-Modal Network with Dynamic Feature Fusion for Emotional Mimicry Intensity Estimation

### v2

@@ Abstract
Emotional Mimicry Intensity (EMI) estimation serves as a critical technology for understanding human social behavior and enhancing human-computer interaction experiences, where the core challenge lies in dynamic correlation modeling and robust fusion of multimodal temporal signals. To address the limitations of existing methods in insufficient exploitation of modal synergistic effects, noise sensitivity, and limited fine-grained alignment capabilities, this paper proposes a dual-stage cross-modal alignment framework. First, we construct vision-text and audio-text contrastive learning networks based on an improved CLIP architecture, achieving preliminary alignment in the feature space through modality-decoupled pre-training. Subsequently, we design a temporal-aware dynamic fusion module that combines Temporal Convolutional Networks (TCN) and gated bidirectional LSTM to respectively capture the macro-evolution patterns of facial expressions and local dynamics of acoustic features. Innovatively, we introduce a quality-guided modality fusion strategy that enables modality compensation under occlusion and noisy scenarios through differentiable weight allocation. Experimental results on the Hume-Vidmimic2 dataset demonstrate that our method achieves an average Pearson correlation coefficient of 0.51 across six emotion dimensions, providing a novel technical pathway for fine-grained emotion analysis in open environments.
**footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
@@ 1 Introduction
This study is based on the EMI track of the 8th ABAW (Affective Behavior Analysis in-the-wild) Competition[20, 19, 14, 17, 26, 18, 23, 24, 15, 16, 21, 11, 12, 22, 13, 25]. The adopted dataset, Hume-Vidmimic2, comprises 15,000 videos totaling over 25 hours of content.Participants in these videos mimic emotional expressions observed in seed videos, with their performance subsequently evaluated across multiple dimensions including ”admiration,” ”amusement,” ”determination,” ”empathic pain,” ”excitement,” and ”joy.”
Through systematic experiments on the Hume-Vidmimic2 dataset, this paper validates the effectiveness of the proposed method. Compared to baseline models, our framework achieves a breakthrough performance with an average Pearson correlation coefficient of 0.51 across six emotional dimensions, demonstrating 200% improvement over the best baseline (0.25).
@@ 3.2 Overall Framework
The proposed multimodal EMI estimation framework incorporates a dual-stage training paradigm and multimodal fusion architecture. The first stage achieves independent semantic alignment between image-text and audio-text pairs through contrastive learning. The second stage constructs cross-modal temporal modeling and dynamic fusion modules while freezing modality encoders. Final emotional mimicry intensity estimation is implemented end-to-end via a multi-layer Transformer encoder.
@@ 3.3 Dual-Stage Cross-Modal Alignment
Stage I: Modality-Text Contrastive Alignment Building upon the CLIP architecture [28], we establish dual-path contrastive learning networks for vision-text and audio-text pairs. For visual modality, given video frame sequence $V=\{v_{t}\}_{t=1}^{T}$, frame-wise features $f_{v}^{t}\in\mathbb{R}^{d_{v}}$ are extracted through a ViT encoder, with segment-wise average pooling yielding video-level global features $\bar{f}_{v}$. Simultaneously, textual modality employs a pretrained BERT encoder to map transcribed speech $S$ into text features $f_{s}\in\mathbb{R}^{d_{s}}$. The vision-text contrastive loss is defined as:
$\mathcal{L}_{v2t}=-\log\frac{\exp(\bar{f}_{v}\cdot f_{s}/\tau)}{\sum_{i=1}^{N}\exp(\bar{f}_{v}\cdot f_{s_{i}}/\tau)}$ (1)
where $\tau$ denotes the temperature coefficient and $N$ the batch size. For audio modality, the CLAP architecture [32] processes acoustic signals $A$ through Wav2Vec2.0 [2] to extract spectral features $f_{a}\in\mathbb{R}^{d_{a}}$, with analogous audio-text contrastive loss $\mathcal{L}_{a2t}$ following Equation (1).
@@ 4.1 Datasets and Experimental Setup
Experiments are conducted on the Hume-Vidmimic2 multimodal affective dataset, which contains 15,000 video samples spanning six emotional dimensions (”Admiration”, ”Amusement”, ”Determination”, ”Empathic Pain”, ”Excitement”, ”Joy”). Each sample’s emotional intensity score, ranging within $[0,1]$, is annotated by professional annotation teams following standardized protocols. The dataset is partitioned into training, validation, and test sets. Statistical details of each split are provided in Table  1.
@@ 5.1 Unimodal Performance Analysis
As shown in Table 2, our enhanced unimodal models demonstrate consistent improvements over baseline implementations across both modalities. For the visual modality, it can be observed that our ViT model pre-trained using contrastive learning demonstrates significantly superior performance compared to the official ViT features. our modified ViT architecture achieves a Pearson correlation coefficient ($\rho_{val}$) of 0.16, representing a 77.8% relative improvement over the baseline’s 0.09. In the acoustic domain, our optimized Wav2Vec2.0 implementation reaches $\rho_{val}$=0.2734, outperforming the baseline (0.24) by 13.9%. This significant performance gain in audio processing suggests our feature enhancement strategy effectively captures subtle paralinguistic cues critical for emotional mimicry estimation.
5.2 Multimodal Fusion Effectiveness
  • The audio-visual pair achieves $\rho_{val}$=0.3746, surpassing visual-textual fusion (0.2834) by 32.1%
  • The audio-textual pair demonstrates strong synergy with $\rho_{val}$=0.4418, highlighting text modality’s complementary role
The progressive performance gains confirm our hypothesis about cross-modal complementarity. Notably, the textual modality contributes disproportionately to system performance, suggesting linguistic and audio context plays a crucial role in disambiguating emotional mimicry patterns. Our dynamic fusion mechanism successfully preserves modality-specific information while enabling effective cross-modal interaction, as evidenced by the super-additive performance improvement in three-modality fusion.
@@ 5.2 Multimodal Fusion Effectiveness
Table 3: The Multimodal results on validation set of the EMI Estimation Challenge.We use our pre-trained encoders for three modalities to conduct ablation experiments.
@@ 6 Conclusion
This paper presents our solution to the Emotional Mimicry Intensity (EMI) estimation challenge in the 8th Affective Behavior Analysis in-the-wild (ABAW) Competition. By introducing a dual-stage training paradigm and quality-aware dynamic modality fusion mechanism, our method achieves enhanced robustness in EMI estimation while maintaining tight three-modality alignment. Systematic validation on the Hume-Vidmimic2 dataset demonstrates that our approach attains an average Pearson correlation coefficient of 0.51 across six emotional dimensions.Provide a new approach for emotional mimicry intensity estimation.

### v3

# Technical Approach for the EMI Challenge in the 8th Affective Behavior Analysis in-the-Wild Competition

@@ Abstract
Emotional Mimicry Intensity (EMI) estimation plays a pivotal role in understanding human social behavior and advancing human-computer interaction. The core challenges lie in dynamic correlation modeling and robust fusion of multimodal temporal signals. To address the limitations of existing methods—insufficient exploitation of cross-modal synergies, sensitivity to noise, and constrained fine-grained alignment capabilities—this paper proposes a dual-stage cross-modal alignment framework. Stage 1 develops vision-text and audio-text contrastive learning networks based on a CLIP architecture, achieving preliminary feature-space alignment through modality-decoupled pre-training. Stage 2 introduces a temporal-aware dynamic fusion module integrating Temporal Convolutional Networks (TCN) and gated bidirectional LSTM to capture macro-evolution patterns of facial expressions and local dynamics of acoustic features, respectively. A novel quality-guided fusion strategy further enables differentiable weight allocation for modality compensation under occlusion and noise. Experiments on the Hume-Vidmimic2 dataset demonstrate superior performance with an average Pearson correlation coefficient of 0.51 across six emotion dimensions on the validate set. Remarkably, our method achieved 0.68 on the test set, securing runner-up in the EMI Challenge Track of the 8th ABAW (Affective Behavior Analysis in the Wild) Competition, offering a novel pathway for fine-grained emotion analysis in open environments.
@@ 1 Introduction
This study is based on the EMI track of the 8th ABAW (Affective Behavior Analysis in-the-wild) Competition[21, 20, 15, 18, 27, 19, 24, 25, 16, 17, 22, 11, 12, 23, 13, 26]. The adopted dataset, Hume-Vidmimic2, comprises 15,000 videos totaling over 25 hours of content.Participants in these videos mimic emotional expressions observed in seed videos, with their performance subsequently evaluated across multiple dimensions including "admiration," "amusement," "determination," "empathic pain," "excitement," and "joy."
Through systematic experiments on the Hume-Vidmimic2 dataset, this paper validates the effectiveness of the proposed method. Compared to baseline models, our framework achieves a breakthrough performance with an average Pearson correlation coefficient of 0.51 across six emotional dimensions, demonstrating 112% improvement over the best baseline (0.24).
@@ 3.1 Multimodal Data Preprocessing
Figure 1: The overview of our proposed method. We first utilize the three encoders with frozen parameters pre-trained in the first stage, the audio, image, and text are then fed into these three encoders to obtain $F_{A}$, $F_{I}$, $F_{T}$. After that, $F_{I}$ and $F_{A}$ will be strengthened by temporal features for the Temporal Feature Enhancer (TFE), the temporally enhanced features are then processed by the Quality-Aware Module (QAM), which outputs weights for each feature. These weights are subsequently applied to the corresponding features, and the weighted features from the three modalities (audio, visual, text) are concatenated and fed into the Transformer Decoder module followed by the Classifier module to generate the final prediction.
@@ 3.2 Overall Framework
The proposed multimodal EMI estimation framework incorporates a dual-stage training paradigm and multimodal fusion architecture. The first stage achieves independent semantic alignment between image-text and audio-text pairs through contrastive learning. The second stage constructs cross-modal temporal modeling and dynamic fusion modules while freezing modality encoders. Final emotional mimicry intensity estimation is implemented end-to-end via a multi-layer Transformer encoder. Figure 1 shows an overview of the proposed framework.
@@ 3.3 Dual-Stage Cross-Modal Alignment
Stage I: Modality-Text Contrastive Alignment Building upon the CLIP architecture [29], we establish dual-path contrastive learning networks for vision-text and audio-text pairs. For image-text modality alignment, we utilize the Aff-Wild2[14] multimodal dataset. This database contains 548 in-the-wild videos (approximately 2.7 million frames) annotated with six basic expressions (anger, disgust, fear, happiness, sadness, surprise), the neutral state, and an "other" compound affective category. Additionally, each frame includes Action Unit (AU) annotations and Valence-Arousal (VA) annotations to capture facial muscle movements and emotional intensity dynamics.
To address the underutilization of fine-grained affective semantics in existing facial expression recognition methods, we propose a cross-modal contrastive learning framework leveraging AffectNet’s multimodal annotations. We first integrate discrete emotion labels (8 basic emotions), continuous Valence-Arousal (VA) values, and Action Unit (AU) codes into hierarchical textual descriptions. For each facial image, we generate structured text prompts following this template:
$\begin{split}\mathcal{T}&=\underbrace{\text{"[Expression Class]"}}_{\text{Primary Label}}\\ &\quad+\underbrace{\text{"with [AU Codes]"}}_{\text{Muscular Activity}}\\ &\quad+\underbrace{\text{"Valence=VA1, Arousal=VA2"}}_{\text{Psychological Dimension}}\end{split}$ (1)
For instance: "Happy (Intensity: High) with AU6 (Cheek Raiser) and AU12 (Lip Corner Puller), Valence=0.83, Arousal=0.65". This multi-granularity text augmentation converts the original sparse annotations into semantically dense text.
Architecturally, we adapt the CLIP framework for cross-modal alignment: A VIT-large visual encoder extracts facial features $f_{v}\in\mathbb{R}^{768}$, while a BERT-base text encoder generates textual embeddings $f_{s}\in\mathbb{R}^{768}$. The contrastive loss incorporates annotation confidence weights into InfoNCE:
$\mathcal{L}_{\text{con}}=-\frac{1}{N}\sum_{i=1}^{N}w_{i}\log\frac{\exp(f_{v}^{(i)}\cdot f_{s}^{(i)}/\tau)}{\sum_{j=1}^{N}\exp(f_{v}^{(i)}\cdot f_{s}^{(j)}/\tau)}$ (2)
where weights $w_{i}$ are determined by VA value standard deviations, with $\tau=0.07$ as temperature.
For audio-text modality alignment, we use both Aff-Wild2 and Hume-Vidmimic2[27] dataset. Specially, for audio modality, the CLAP architecture [33] processes acoustic signals through Wav2Vec2.0 [2] ,which extract audio features $f_{a}\in\mathbb{R}^{d_{a}}$. The remaining parts follow the same image-text alignment process as described above.
@@ 4.1 Datasets and Experimental Setup
Experiments are conducted on the Hume-Vidmimic2[27] multimodal affective dataset, which contains 15,000 video samples spanning six emotional dimensions ("Admiration", "Amusement", "Determination", "Empathic Pain", "Excitement", "Joy"). Each sample’s emotional intensity score, ranging within $[0,1]$, is annotated by professional annotation teams following standardized protocols. The dataset is partitioned into training, validation, and test sets. Statistical details of each split are provided in Table  1.
@@ 5.1 Unimodal Performance Analysis
As shown in Table 2, our enhanced unimodal models demonstrate consistent improvements over baseline implementations across both modalities. For the visual modality, it can be observed that our ViT model pre-trained using contrastive learning demonstrates significantly superior performance compared to the official ViT features. our modified ViT architecture achieves a Pearson correlation coefficient ($\rho_{val}$) of 0.1481, representing a 70% relative improvement over the baseline’s 0.0873. In the acoustic domain, our optimized Wav2Vec2.0 implementation reaches $\rho_{val}$=0.2734, outperforming the baseline (0.2405) by 13.9%. This significant performance gain in audio processing suggests our feature enhancement strategy effectively captures subtle paralinguistic cues critical for emotional mimicry estimation.
@@
5.2 Multimodal Fusion Effectiveness
As shown in Table 3, the Quality-Aware Module (QAM) demonstrates critical efficacy in multimodal emotion analysis. The baseline model (VIT+Wav2Vec+BERT) achieves a Pearson correlation coefficient $\rho_{val}$ of 35.05%, while introducing QAM alone boosts performance to 48.81% (+39.3% relative improvement).This significant performance improvement stems from QAM’s selective feature activation mechanism. When the video modality is corrupted by noise, QAM dynamically shifts fusion weights toward more stable audio and text modalities.
Experimental Combination TFE QAM Mean $\rho_{val}$
VIT+Wav2Vec+BERT 0.3505
+TFE 0.4219
+QAM 0.4881
+TFE + QAM 0.5126
Table 3: Module Ablation Experiment Results on $\rho_{val}$.The ’VIT+Wav2Vec+BERT’ combination serves as our baseline.
5.3 Multimodal Fusion Effectiveness
  • The audio-visual pair achieves $\rho_{val}$=0.4846, surpassing visual-textual fusion (0.2834) a lot.This indicates that audio data contributes disproportionately more than textual data.
  • The audio-textual pair demonstrates strong synergy with $\rho_{val}$=0.4918, highlighting text modality’s complementary role
@@ 5.3 Multimodal Fusion Effectiveness
Table 4: The Multimodal results on validation set of the EMI Estimation Challenge.We use our pre-trained encoders for three modalities with both TFE and QAM to conduct ablation experiments.
The progressive performance gains confirm our hypothesis about cross-modal complementarity. Notably, the textual and audio modality contribute disproportionately to system performance, suggesting linguistic and audio context play a crucial role in disambiguating emotional mimicry patterns. Our dynamic fusion mechanism successfully preserves modality-specific information while enabling effective cross-modal interaction, as evidenced by the super-additive performance improvement in three-modal fusion.
@@ 6 Conclusion
This paper presents our solution to the Emotional Mimicry Intensity (EMI) estimation challenge in the 8th Affective Behavior Analysis in-the-wild (ABAW) Competition. By introducing a dual-stage training paradigm and quality-aware dynamic modality fusion mechanism, our method achieves enhanced robustness in EMI estimation while maintaining tight three-modality alignment. Systematic validation on the Hume-Vidmimic2 dataset demonstrates that our approach attains an average Pearson correlation coefficient of 0.51 across six emotional dimensions.Moreover, our method achieved an average Pearson correlation coefficient of 0.68 on the test set, secured the runner-up position in the 8th ABAW (Affective Behavior Analysis in the Wild) Competition, demonstrating excellent performance in emotion recognition tasks under uncontrolled real-world conditions. Provide a new approach for emotional mimicry intensity estimation.